OpenStudy (anonymous):

Which of the following is equivalent to \[\sum_{n=3}^{5} 2n+1\] ?

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

27 18 36 30

OpenStudy (luigi0210):

Plug them in and add: 7+9+11

OpenStudy (agent0smith):

That means you need to add up 2n+1, starting from n=3 and stopping at n=5:\[\Large \sum_{n=3}^{5} 2n+1 = (2*3+1) +(2*4+1) +(2*5+1)\]

OpenStudy (anonymous):

27?

OpenStudy (anonymous):

thank you both!